OpenStudy (lgbasallote):

PLEASE HELP!!! IT'S 3 AM HERE AND I NEED THIS HOMEWORK EXPLAINED!!! OpenStudy user s.k has 343 fans less than a64. If s.k gets 3 fans a week while a64 gets 2 fans a week, how long will it take s.k to get twice the fans a64 has?

OpenStudy (saifoo.khan):

LOL. nice thought.

OpenStudy (saifoo.khan):

s.k. fans = x a64 fans = y y = 343 + x i get till here..

OpenStudy (lgbasallote):

im thinking y + 2a = x + 3a?

OpenStudy (akshay_budhkar):

After z weeks the number of fans of s.k= x+3z and a64= y+2z

OpenStudy (lgbasallote):

no wait..that's for them to equal..

OpenStudy (lgbasallote):

if s.k. will double then 2(y + 2z) = x + 3z?

OpenStudy (akshay_budhkar):

now we want them to be twice, so we want x+3z=2(y+2z)

ganeshie8 (ganeshie8):

y1 = 3/7 x y2 = 2/7 x + 343

OpenStudy (lgbasallote):

x + 3z = 2y + 4z x - 2y = z what does that mean?

OpenStudy (lgbasallote):

x - 2(343 + x)?

ganeshie8 (ganeshie8):

y1 = 3/7 x y2 = 2/7 x + 343 y1 = 2y2 solve for x

OpenStudy (lgbasallote):

how did you get those @ganeshie8 ?

ganeshie8 (ganeshie8):

y1: number of fans of sk y2 : number of fans of a64 y1 is increasing his fans at a rate of 3 per 7 days right so slope of y1 is 3/7 put y1 equation in slope intercept form

OpenStudy (akshay_budhkar):

yea agreed^ why did i introduce a new variable O_O?

OpenStudy (saifoo.khan):

lol ^

ganeshie8 (ganeshie8):

to start with, a64 has 343 as y intercept.. hope you get the idea..

OpenStudy (lgbasallote):

3/7 x = 2(2/7x +343) 3/7 x - 4/7 x = 868 -1/7 x = 868??? this seems weird

OpenStudy (lgbasallote):

uhm it's 686 lol

OpenStudy (lgbasallote):

but still negative so weird o.O

OpenStudy (saifoo.khan):

So sk should start asking people to unfan him? :O

OpenStudy (lgbasallote):

lol =)))))

OpenStudy (saifoo.khan):

Sounds like a plan. NOT.

OpenStudy (lgbasallote):

s.k should put up a teenage asian girl...

OpenStudy (lgbasallote):

pic

OpenStudy (saifoo.khan):

Ok, this sounds like a better plan. :D

OpenStudy (saifoo.khan):

what is this guy typing from a hour?

OpenStudy (lgbasallote):

even with values im getting negative lol

ganeshie8 (ganeshie8):

me.. ? im not typing anythign... just watching hindi movie i thoutht lgba got the answer... ?

OpenStudy (saifoo.khan):

No, i was talking about lgba.

OpenStudy (lgbasallote):

no i didnt get the answer..im getting negatives

OpenStudy (lgbasallote):

the time when they equal is just easy...could it be that it's impossible for s.k to get double the fans of a64?? o.O

OpenStudy (saifoo.khan):

Possibility..

ganeshie8 (ganeshie8):

yeah that can happen in reality... but equations.. we should get +ve numb er of days

OpenStudy (lgbasallote):

let's try solving for equal first... x + 3z = y + 2z z = y - x <---this is the time when they become equal

OpenStudy (lgbasallote):

assuming they are already equal let a be their values a + 3n = 2(a + 2n) a + 3n = 2a + 4n a - 2a = n <---the time when they double

OpenStudy (lgbasallote):

so if z + n = -a + a = 0??? o.O

OpenStudy (saifoo.khan):

LOL.

OpenStudy (akshay_budhkar):

equal? u want it twice..

OpenStudy (lgbasallote):

wth am i doing wrong =_=

OpenStudy (lgbasallote):

well it didnt work lol..i just tried

OpenStudy (lgbasallote):

why am i getting negative @Akshay_Budhkar ? any ideas>

OpenStudy (akshay_budhkar):

i am working on it gimme a moment

OpenStudy (lgbasallote):

ugh i gtg...will be back in an hour... if any one of you got some ideas pls feel free to post it here

ganeshie8 (ganeshie8):

ok me too confused..

OpenStudy (lgbasallote):

i dont think i did anything wrong in the solution :/

OpenStudy (akshay_budhkar):

are we missing an equation??

OpenStudy (akshay_budhkar):

yea we need another equation, the answer will depend on the initial values of a and b!! try taking two diff sets of the numbers and solving

OpenStudy (akshay_budhkar):

@saifoo.khan ^

OpenStudy (anonymous):

I don't think you can solve this until you have another equation. We have 3 variables (a= s.k fans, b=a64 fans, and w=weeks) but only two equations, so it is unsolvable. right?

OpenStudy (lgbasallote):

well x and y dont really matter. as long as you get a numerical value...but let's say s.k has 914 fans and a64 was 1257..it still becomes negative

OpenStudy (lgbasallote):

@ganeshie8 still nothing? even if there are values?

ganeshie8 (ganeshie8):

343 + 3t = 2{ 2t} => t = 343 weeks in 343 weeks sk fans will become twice that of a64. i am still lost in straight lines.. i want to see you or someone else explain the answer @lgbasallote @shubhamsrg @abstracted @Akshay_Budhkar

ganeshie8 (ganeshie8):

@mathslover

OpenStudy (lgbasallote):

how did you get 343 again?

OpenStudy (blockcolder):

You can just say that every week, s.k's lead will be reduced by one so that after 343 weeks, s.k's lead will be zero.

OpenStudy (lgbasallote):

wait what?

OpenStudy (akshay_budhkar):

@ganeshie8 the straight line stuff wont work, the intercept is wrong, you dont know the intercept. @lgba the question is incomplete...

OpenStudy (saifoo.khan):

@lgbasallote , come up with the complete question next time! -_- i want to find a way out.

ganeshie8 (ganeshie8):

@Akshay_Budhkar ,straight line logic may not tell the answer straight away.. but it works. @saifoo.khan , the question @lgbasallote posted is complete. answer is : sk can never achieve his goal. because with a 3t rate he can never reach 2(2t) value !! 3t < 2 (2t + 343) 3t < 4t + ....

OpenStudy (akshay_budhkar):

@ganeshie8 it is pretty illogical... he can achieve his goal!! He is going at a faster rate man

OpenStudy (akshay_budhkar):

lol take real life example of 0 fans and 343 fans and solve, it works

ganeshie8 (ganeshie8):

for positive values of t, 3t <= 2(2t + 300 ) u agree... ?

ganeshie8 (ganeshie8):

@Akshay_Budhkar

OpenStudy (lgbasallote):

but how could it not be possible?? o.O i mean soon s.k is gonna overtake a64 then after that s.k will always be leading there should be a time when s.k will get twice :O

OpenStudy (lgbasallote):

let's say both of them start at 0... 1st week - s.k is 3... a64 is 2 2nd week - 6 - 4 3rd week - 9 - 6 4th week - 12 - 8 5th week - 15 - 10 6th week - 18 -12 see how s.k's lead is gettingbigger...that should double sometime

OpenStudy (anonymous):

Let \(x\) be a64's number of fans. Let \(y\) be s.k.'s number of fans. We have \(y(x; w=0)=x_i-343.\) which is the initial value for \(y\) based on the initial value of \(x.\) Based on how they gain fans, we have: \(y(x,w)=x_i-343+3w.\) We want to find the point at which \(y(x,w)=2x(w)\). Note that \(x(w)=x_i+2w.\) Thus we have, \(x_i-343+3w=2x_i+4w.\) Let's let \(x_i\) be a constant. So, we have: \(-x_i-343=w\) We assume that \(x_i \in \mathbb{Z}^+.\) This tells us that \(y(x,w)\) will never have twice as many fans as \(x(w)\) if the functions behave as they have been defined. It actually says that we would have to go "backward" to reach a point where \(y(x,w)\) had twice as many fans as \(x(w)\). Also, you contradicted yourself, lgba: "let's say both of them start at 0..." This contradicts the premise that they would start in relation to one another. (i.e. \(y(x;w=0)=x_i-343\) and \(x(w)=x_i+2w.\)) Now let's suppose this: Both \(x\) and \(y\) start at some value \(x_0\) and \(y_0\) respectively. Their growth functions are \(x(w)=x_0+2w\) and \(y(w)=y_0+3w.\) If we want to find the point at which \(y(w)=2x(w),\) we have \(y_0+3w=2x_0+4w\) which is \(w=y_0-2x_0.\) You can see how this makes no sense in your original presumptions. We can continue this logic further and deduce that this problem makes no sense if \(x_0 > \frac{y_0}{2}.\) What I find most interesting is that while there is no point in future time that \(y(w)=2x(w),\) there is a point where \(y(w)=x(w).\) This happens to be \(w=x_0-y_0.\) It is this fact that makes one speculate that there is a point where \(y(w)=2x(w)\), but there is no such point in future time.

OpenStudy (anonymous):

Oh, cool! I also figured out something else: \[ \begin{align} f(x)&=m_fx+b_f\\ g(x)&=m_gx+b_g\\ \lim_{x \to \infty }\frac{f(x)}{g(x)}&=\frac{m_f}{m_g} \end{align} \] This explains everything.

OpenStudy (anonymous):

(Note: that infinity should be \(+\infty.\))

ganeshie8 (ganeshie8):

@lgbasallote let's say both of them start at 0... 1st week - s.k is 3... a64 is 2 2nd week - 6 - 4 3rd week - 9 - 6 4th week - 12 - 8 5th week - 15 - 10 6th week - 18 -12 . . . 100th week - 300 - 200 1000th week - 3000 - 2000 10000th week - 30000 - 20000 lol both must have died by now... the trick is "a 3t rate can never increase its value to 2(2t)"

OpenStudy (anonymous):

@ganeshie8, I just said that. You essentially copied my work.

ganeshie8 (ganeshie8):

lol.. i copied your work :)

OpenStudy (lgbasallote):

but...but... T_T sorry @saifoo.khan seems amistre is unbeatable

OpenStudy (lgbasallote):

but...but... T_T sorry @saifoo.khan seems amistre is unbeatable

OpenStudy (anonymous):

@lgbasallote, you can get above amistre. You simply cannot get to where you are double him.

OpenStudy (lgbasallote):

yeah...what i meant lol

OpenStudy (saifoo.khan):

@lgbasallote = Good for nothing!!! =__= Jk